# Passive Snapshot Coded Aperture Dual-Pixel RGB-D Imaging

### v1

@@ 1 Introduction
Figure 1: (A). We propose CADS - an imaging approach that leverages DP sensors and coded aperture masks for passive snapshot 3D imaging. (B) Coded DP sensing improves on naive dual-pixel sensing for simultaneous depth estimation and deblurring. Our CADS imaging prototype can recover accurate depth maps and all-in-focus images in (C) DSLR photography settings, as well as for (D) 3D endoscopy and dermoscopy.
Extracting 3D information from 2D images has widespread applications in several domains such as microscopy, medical imaging, and autonomous driving. Conventional methods like passive stereo, structured illumination [5, 24], photometric stereo [3], and time-of-flight sensing do an excellent job estimating depth maps from 2D images for a wide range of computer vision tasks. However, these 3D imaging methods have a bulkier form factor due to the use of multiple cameras/illumination sources, and their estimation accuracy suffers when they are used in systems with size restrictions.
Monocular depth estimation has the potential to overcome these form factor restrictions. One popular monocular depth estimation strategy is coded aperture imaging in which the aperture of the lens is coded either using an amplitude [16, 25] or a phase mask [31]. However, the quality of the depth maps and all-in-focus images obtained using these coded-aperture systems is still subpar compared to traditional methods like stereo or time-of-flight sensing. Recently, dual-pixel sensors have emerged as accurate monocular depth sensing technology that overcomes these drawbacks due to their unique sensor design. These sensors are commonly available in smartphones and DSLR cameras.
@@
To improve the depth-of-field of dual-pixel cameras while still maintaining their ability to capture depth information, we introduce a novel sensing strategy called CADS - Coded Aperture Dual-pixel Sensing. CADS uses a learned amplitude mask (code) in the aperture plane of a camera equipped with a DP sensor to improve the conditioning of the DP defocus blur while maintaining the disparity estimation ability from the two views. We use a two-stage, end-to-end learning framework to simultaneously learn the optimal coded mask pattern and the optimal neural network weights that can produce an accurate depth map and deblurred all-in-focus image estimates from dual-pixel measurements for a wide range of aperture settings, see Fig. 2.
Compared to conventional DP sensing (with no coded aperture), the learned coded dual-pixel sensing system provides a better AIF performance, indicated through a consistent $\geq 1.5$dB PSNR improvement in the quality of AIF images and a 5-6% improvement in the quality of depth maps for a wide range of aperture settings, allowing a better disparity and AIF quality trade-off. We verify the efficacy of CADS through various simulations and real-world experiments. CADS provides a mechanism for high-fidelity depth and intensity imaging in a small form factor. We demonstrate its benefits in two relevant real-world applications - 3D endoscopy and EDOF dermoscopy. For both applications, we show promising real-world experimental results using proof-of-concept prototypes built in the lab, see Fig. 1.
In summary, our contributions include:
  • We propose coded aperture dual-pixel sensing called CADS that significantly improves the trade-off between quality of disparity and AIF estimates from dual-pixel sensors.
  • At the core of our method is the proposed differentiable coded dual-pixel renderer that allows end-to-end learning of the coded aperture mask-specific dual-pixel geometry.
@@ 3.1 Coded dual-pixel image formation model
However, this disparity comes with an undesirable effect of defocus blur. In a DP imaging system, the defocus blur size is directly proportional to the disparity. To have good accurate depth measurements, one needs to have a large disparity range between the closest and the farthest scene depths, but this comes at the cost of highly defocused, blurry images, leading to poor all-in-focus image estimation.
@@
When a coded (amplitude) mask $M$ is added to an imaging system/camera with DP sensor, the coded DP PSFs can be expressed as a Hadamard multiplication of the naive DP PSFs with the coded mask. This is illustrated in Fig. 3. More formally, the left and right DP PSFs ($h^{L,C}_{z}$ and $h^{R,C}_{z}$ respectively) can be expressed as
@@ 3.2.1 Coded dual-pixel image simulator
Mask-to-PSF generation. In our CADS framework, we parameterize our mask pattern $M$ with a continuous-valued map $\theta_{c}(x,y)$ as done in [25], where $M_{\theta_{c}}(x,y)=\sigma(\alpha*\theta_{c}(x,y))$, where $\sigma(.)$ is the sigmoid function and $\alpha$ is the temperature factor controlling mask transparency. For realistic modeling of the naive (no-code) DP PSFs $h^{L}_{z}$, $h^{R}_{z}$, we use the formulation provided in Abuolaim et al. [2]. We generate left and right naive DP PSFs for $21$ depth planes, with signed circle-of-confusion (CoC) sizes ranging from $-40$ pixels to $+40$ pixels. More details regarding this can be found in the supplementary. Given $M_{\theta_{c}}$, $h^{L}_{z}$, $h^{R}_{z}$, we use Eq. 1 to generate the coded left, right DP PSF z-stack i.e. $h^{L,C}_{z}$ and $h^{R,C}_{z}$ for all the $21$ depth planes.
Coded DP image rendering. We adopt a multi-plane image (MPI) representation of the scene, where the scene is divided into discrete depth planes. Given $h^{L,C}_{z}$ and $h^{R,C}_{z}$, the coded DP left, right images $I_{L}$, $I_{R}$ of a 3D scene can be expressed as follows -
@@
where $s_{z}$ is the scene intensity at depth $z$, and *** is the 2D convolution operator. To remove artifacts at the edges of the MPI depth layers, we adopt the differentiable occlusion-aware modifications to Eqn. 2 from Ikoma et al. [11], along with a minor modification. See supplementary for details.
@@ 3.2.2 CADNet for defocus and AIF estimation
CADNet takes as input the blurry measurements and outputs - (1) a normalized defocus map $\hat{D}_{N}$ ranging from -1 to 1, and (2) the deblurred all-in-focus (AIF) image $\hat{Y}$ of the scene. The normalized defocus map $\hat{D}_{N}$ is converted to actual physical defocus blur size by scaling it as $\hat{D}=\hat{D}_{N}D_{max}/p$, where $p$ is the pixel pitch of the DP sensor, and $D_{max}$ being the maximum blur size (in px) that CADNet is trained to handle. This defocus value is related to the imaging depth as follows [8]
@@
$\begin{split}\mathcal{L}=\mathcal{\beta}_{1}\mathcal{L}_{L1}(\hat{Y},Y_{gt})+\beta_{2}\mathcal{L}_{L1}(g(\hat{Y}),g(Y_{gt}))+\\ \beta_{3}\mathcal{L}_{L1}(\hat{D},D_{gt})+\beta_{4}\mathcal{L}_{L1}(g(\hat{D}),g(D_{gt}))+\\ \beta_{5}\mathcal{L}_{mask},\end{split}$ (4)
where $\mathcal{L}_{L1}$ is the $L_{1}$ loss, $g(.)$ is 2D gradient operator, and $Y_{gt}$ and $D_{gt}$ are groundtruth AIF and defocus maps. The last term is a mask pattern regularizer $\mathcal{L}_{mask}=\text{Relu}\left(0.5-\sum_{x,y}M_{\theta_{C}}/\sum_{x,y}M_{open}\right)$, where $M_{open}$ is the open aperture mask pattern. We set $(\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5})=(1,0.5,1,0.5,1,1000)$, with $\beta_{5}$ being large to enforce a strict transmission constraint ($\geq 50\%$ light efficiency).
@@ 3.2.3 Implementation details
For training the CADS framework, we used simulated data generated using the forward model discussed in Section 3.2 from the FlyingThings3D dataset[17]. 20k scenes are split $80\%-20\%$ into training and validation sets. We rescale the disparity maps for each scene to convert them to depth maps with values lying within $\mathcal{D}_{T}$ = [$32$ mm, $76$ mm]. During training our coded DP simulator, we fix our camera parameters to be $f=4.38$ mm, $L=f/1.73$, $g=45$ mm, and our depth range corresponding to CoC sizes of $-40$ to $40$ pixels. Based on Eq. 3, this amounts to a depth range of $\mathcal{D}_{T}$. During testing on simulated data, we use the following parameters: focal length $f=50$ mm, aperture $L=f/4$, in-focus distance $g=40$ cm, with a depth range $\mathcal{D}_{V}$ = [$32$ cm, $52$ cm]. The signed blur sizes correspond to -40 to 40 pixels, assuming a pixel pitch of 10.72 $\mu$m. We learn a coded mask pattern with light efficiency $>=50\%$. To account for light loss due to coded mask, we add appropriate heteroscedastic noise [6] to simulated captures. We train the end-to-end framework using the Adam optimizer. We initialize the coded aperture mask with an open aperture pattern of size 21x21 pixels. We initialize $\alpha_{0}$ as 0 and increase it using the schedule $\alpha_{t}=\alpha_{0}+\frac{t}{8000}$, where $t$ is the total number of iterations. For CADNet weights, we set the learning rate to be 3× $3\,\times\,10^{-5}$, and for learning the mask pattern, we set the learning rate to 3× $3\,\times\,10^{-4}$. We use cosine decay scheduling. More details are in the supplementary.
@@ 4.1 Comparison with naive standard and dual pixel sensing
As a result of end-to-end training, we learn an optimal coded mask pattern that better conditions the coded DP blurry measurements, thus helping in good depth and AIF recovery. The learned mask and the corresponding left, right coded DP PSFs as a result of our end-to-end training have been depicted in Figs. 2, 3. In this section, we show that our learned coded-aperture DP sensing is in fact an improvement over naive monocular standard and dual-pixel sensing in terms of both AIF and depth estimation quality. We use PSNR, SSIM, and LPIPS measures[34] for the evaluation of AIF quality. We use Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and depth estimation accuracy within a threshold margin ($\delta<1.05^{1}$) for the evaluation of depth prediction quality.
@@
Figure 5: Comparison with naive standard and DP - real data. From left to right we show AIF and depth predictions by naive standard-pixel, naive dual-pixel and CADS, respectively, along with the ground truth. CADS gives the best depth and AIF predictions. GT for AIF is obtained by capturing the scene with a $f/22$ aperture. A coarse GT depth map is obtained using the Intel RealSense sensor. Zoom in for the best result.
@@ 4.2 Performance of CADS for various aperture sizes
CADS offers a better trade-off between depth and AIF quality compared to a naive dual-pixel sensor for various aperture sizes as shown in Fig. 1. To establish this, we evaluate the performance of CADS for various aperture sizes ranging from $f/4$ to $f/10$. We test this out on the same FlyingThings3D scenes. Apart from comparing CADS against naive dual-pixel sensing (Naive DP), we also compare it against - code and no-code variants of standard pixel sensor, referred to as Coded and Naive SP respectively. Table 2 shows depth estimation and AIF deblurring results on our simulated FlyingThings3D data, where the imaging system parameters are the same as mentioned before, with objects being placed in the [32 cm, 53 cm] range. With decreasing aperture size (increasing f-number), AIF prediction accuracy increases while depth prediction gets worse. Our proposed E2E CADS system consistently shows the best performance across all apertures and improves the defocus-disparity trade-off.
@@ 4.3 Comparison with existing dual-pixel works
We compare the performance of CADS with existing learning-based dual-pixel sensing works, namely DPDNet[2], DDDNet[20], Xin et al. [32], Punnapurath et al. [22] and Kim et al. [15]. Since existing works were developed for naive (no-code) dual-pixel sensors, we use the naive dual-pixel blur to render captures to evaluate these methods. For evaluating CADS, we render using the coded-aperture DP PSF. We use the validation subset of 2000 images from the FlyingThings3D dataset to evaluate. The results are shown in Table 3. Most of the existing works estimate disparity which is related to defocus map by an unknown scale. For evaluation, we use the affine invariant version of MAE (AI(1)) for disparity estimation quality [8] and PSNR for AIF quality. Note that we use the normalized defocus map output by CADNet for this comparison instead of converting it to a depth map. Works that were designed for unidirectional disparity [20, 32] show poor results when tested on simulated scenes with bidirectional disparity between the DP images. Our proposed CADS outperforms existing methods on the simulated FlyingThings3D dataset. Moreover, for the same naive DP PSF, our CADNet, referred to as Naive DP in Table 3, also outperforms existing methods. We also compare against these methods on rendered images from NYUv2 dataset[27] and report the performance in the supplementary.
@@
Table 3: Comparison with existing DP methods. CADS offers the best AIF and disparity estimation quality on our simulated dataset. Red highlights best, orange highlights second best

### v2

@@ 1 Introduction
Figure 1: (A). We propose CADS - an imaging approach that leverages DP sensors and coded aperture masks for passive snapshot 3D imaging. (B) Coded DP sensing improves on naive dual-pixel sensing for simultaneous depth estimation and deblurring. Our CADS imaging prototype can recover accurate depth maps and all-in-focus images in (C) DSLR photography settings, as well as for (D) 3D endoscopy and dermoscopy. Read more on our website https://shadowfax11.github.io/cads/.
Extracting 3D information from 2D images has widespread applications in several domains such as microscopy, medical imaging, and autonomous driving. Conventional methods like passive stereo, structured illumination [6, 25], photometric stereo [4], and time-of-flight sensing do an excellent job estimating depth maps from 2D images for several computer vision tasks. However, these 3D imaging methods have a bulkier form factor due to the use of multiple cameras/illumination sources, and their estimation accuracy suffers when they are used in systems with size restrictions.
Monocular depth estimation has the potential to overcome these form factor restrictions. A popular monocular depth sensing strategy is coded aperture imaging, in which the aperture of the lens is coded either using an amplitude [17, 26] or a phase mask [32]. However, the quality of the depth maps and all-in-focus images obtained using these coded-aperture systems is still subpar compared to traditional methods like stereo or time-of-flight sensing. Recently, dual-pixel (DP) sensors have emerged as an accurate monocular depth sensing technology that overcomes these drawbacks due to their unique sensor design. These sensors are commonly found in smartphones and DSLR cameras.
@@
To improve the depth-of-field of dual-pixel cameras while still maintaining their ability to capture depth information, we introduce a novel sensing strategy called CADS - Coded Aperture Dual-pixel Sensing. CADS uses a learned amplitude mask (code) in the aperture plane of a camera equipped with a DP sensor to improve the conditioning of the DP defocus blur while maintaining the disparity estimation ability from the two views. We use a two-stage, end-to-end learning framework to simultaneously learn the optimal coded mask pattern and the optimal neural network weights that can produce an accurate depth map and deblurred all-in-focus image estimates from dual-pixel measurements for a wide range of aperture settings (see Fig. 2).
Compared to conventional DP sensing (with no coded aperture), the learned coded dual-pixel sensing system provides a better AIF performance, indicated through a consistent $\geq 1.5$dB PSNR improvement in the quality of AIF images and a 5-6% improvement in the quality of depth maps for a wide range of aperture settings, allowing a better disparity and AIF quality trade-off. We verify the efficacy of CADS through various simulations and real-world experiments. CADS provides a mechanism for high-fidelity depth and intensity imaging in a small form factor. We demonstrate its benefits in two relevant real-world applications - 3D endoscopy and extended depth-of-field (EDOF) dermoscopy. For both, we demonstrate promising real-world experiments using proof-of-concept prototypes built in the lab, see Fig. 1. In summary, our contributions are:
  • We propose coded aperture dual-pixel sensing, called CADS, that significantly improves the trade-off between quality of disparity and AIF estimates from dual-pixel sensors.
  • At the core of our method is the proposed differentiable coded dual-pixel sensing model that allows end-to-end learning of the coded aperture mask specific to the dual-pixel sensor geometry.
@@ 3.1 Coded dual-pixel image formation model
However, this disparity comes with an undesirable defocus blur effect of defocus blur. In a DP imaging system, the defocus blur size is directly proportional to the disparity. To have good, accurate depth measurements, one needs to have a large disparity range between the closest and the farthest scene depths, but this comes at the cost of highly defocused, blurry images, leading to poor AIF image estimation.
@@
When a coded (amplitude) mask $M$ is added to an imaging system/camera with DP sensor, the coded DP PSFs can be expressed as a Hadamard (element-wise) multiplication of the naive DP PSFs with the coded mask. This is illustrated in Fig. 3. More formally, the left and right DP PSFs ($h^{L,C}_{z}$ and $h^{R,C}_{z}$ respectively) can be expressed as
@@ 3.2.1 Coded dual-pixel image simulator
Mask-to-PSF generation. In our CADS framework, we parameterize our mask pattern $M$ with a continuous-valued map $\theta_{C}(x,y)$ as done in [26], where $M_{\theta_{C}}(x,y)=\sigma(\alpha*\theta_{C}(x,y))$, where $\sigma(.)$ is the sigmoid function. $\alpha$ is a scalar temperature factor that controls mask transparency for smooth learning [^1] We gradually increase $\alpha$ over iterations to ensure the mask is learnt smoothly to be binary (0 or 1). See Supplementary for more details.. For realistic modeling of the naive (no-code) DP PSFs $h^{L}_{z}$, $h^{R}_{z}$, we use the formulation provided in Abuolaim et al. [2]. We generate left and right naive DP PSFs for $21$ depth planes, with signed blur sizes ranging from $-40$ pixels to $+40$ pixels. More details regarding this can be found in the supplementary. Given $M_{\theta_{C}}$, $h^{L}_{z}$, $h^{R}_{z}$, we use Eq. 1 to generate the coded left, right DP PSF z-stack i.e. $h^{L,C}_{z}$ and $h^{R,C}_{z}$ for all the $21$ depth planes.
Coded DP image rendering. We adopt a multi-plane image (MPI) representation of the scene, where the 3D scene is modelled as several discrete depth planes. Given $h^{L,C}_{z}$ and $h^{R,C}_{z}$, the coded DP left, right images $I_{L}$, $I_{R}$ of a 3D scene can be expressed as follows -
@@
where $s_{z}$ is the scene intensity at depth $z$, and * is the 2D convolution operator. To remove artifacts at the edges of the MPI depth layers, we adopt the differentiable occlusion-aware modifications to Eqn. 8 from Ikoma et al. [12], along with a minor modification. See supplementary for details.
@@ 3.2.2 CADNet for defocus and AIF estimation
CADNet takes as input the blurry measurements and outputs - (1) a normalized defocus map $\hat{D}_{N}$, ranging from -1 to +1, and (2) the deblurred all-in-focus (AIF) image $\hat{Y}$ of the scene. The normalized defocus map $\hat{D}_{N}$ is converted to actual physical defocus blur size by scaling it as $\hat{D}=\hat{D}_{N}D_{max}/p$, where $p$ is the pixel pitch of the DP sensor, and $D_{max}$ being the maximum blur size (in px) that CADNet is trained to handle. This defocus value is related to the imaging depth as follows [9]
@@
$\mathcal{L}$ $=\mathcal{L}_{AIF}+\mathcal{L}_{defocus}+\mathcal{L}_{mask},$ (4)
$\mathcal{L}_{AIF}$ $=\beta_{1}\lVert\hat{Y}-Y_{gt}\rVert_{1}+\beta_{2}\lVert\nabla\hat{Y}-\nabla Y_{gt}\rVert_{1},$ (5)
$\mathcal{L}_{defocus}$ $=\beta_{3}\lVert\hat{D}-D_{gt}\rVert_{1}+\beta_{4}\lVert\nabla\hat{D}-\nabla D_{gt}\rVert_{1},$ (6)
$\mathcal{L}_{mask}$ $=\beta_{5}\text{Relu}\left(0.5-\frac{\sum_{x,y}M_{\theta_{C}}}{\sum_{x,y}M_{open}}\right)$ (7)
where .$\lVert.\rVert_{1}$ is the $L_{1}$ loss, (.$\nabla$(.)∇ ( . ) is 2D gradient operator. $\hat{Y}$ and $\hat{D}$ are predicted AIF and defocus maps, while $Y_{gt}$ and $D_{gt}$ are groundtruth AIF and defocus maps. The last term is a mask pattern regularizer, where $M_{open}$ is the open aperture mask pattern. We set $(\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5})=(1,0.5,1,0.5,1,10^{3})$, with $\beta_{5}>>1$ to enforce a strict transmission constraint ($\geq 50\%$ light efficiency).
@@ 3.2.3 Implementation details
For training the CADS framework, we used simulated data generated using the forward model (discussed in Section 3.2) on FlyingThings3D dataset scenes[18]. 20k scenes are split $80\%-20\%$ into training and validation sets. We rescale the disparity maps for each scene to convert them to depth maps having range $\mathcal{D}_{T}$ = [$32$ mm, $76$ mm]. During training of our coded DP simulator, camera parameters were set at $f=4.38$ mm, $L=f/1.73$, $g=45$ mm, ensuring that depth range $\mathcal{D}_{T}$ corresponded to blur sizes of $\leq 40$ pixels (based on Eq. 3). During testing on simulated data, we set camera parameters to match our real-world setting with focal length $f=50$ mm, aperture $L=f/4$, in-focus distance $g=40$ cm, with a depth range $\mathcal{D}_{V}$ = [$32$ cm, $52$ cm]. The signed blur sizes for $\mathcal{D}_{V}$ again corresponded to $\leq 40$ pixels (pixel pitch, $p=10.72\mu m$). Using the loss function described in Eqn. 4, we perform end-to-end training, learning weights for CADNet and learning a coded mask pattern with light efficiency $>=50\%$. To account for light loss due to coded mask, we add appropriate heteroscedastic noise [7] to simulated captures. We train the end-to-end framework using the Adam optimizer. We initialize the coded aperture mask with an open aperture pattern of size 21x21 pixels. More details are in the supplementary.
@@ 4.1 Comparison with naive standard and dual pixel sensing
As a result of end-to-end training, we learn an optimal coded mask pattern that better conditions the coded DP blurry measurements, thus helping in good depth and AIF recovery. The learned mask and the corresponding left, right coded DP PSFs as a result of our end-to-end training have been depicted in Figs. 2, 3. We refer the readers to the Supplementary for further illustrations and explanations. In this section, we show that our learned coded-aperture DP sensing is in fact an improvement over naive monocular standard and dual-pixel sensing in terms of both AIF and depth estimation quality. We use PSNR, SSIM, and LPIPS measures[35] for the evaluation of AIF quality. We use Root Mean Square Error (RMSE), Mean Absolute Error (MAE), and depth estimation accuracy within a threshold margin ($\delta^{1}<1.05^{1}$) for the evaluation of depth prediction quality (see Supplementary for definitions).
@@
Figure 5: Comparison with naive standard and DP - real data. From left to right we show AIF and depth predictions by naive standard-pixel, naive dual-pixel and CADS, respectively, along with ground truth. CADS gives the best depth and AIF predictions. GT for AIF is obtained by capturing the scene with a $f/22$ aperture. A coarse GT depth map is obtained using the Intel RealSense sensor. Zoom in for best viewing.
@@ 4.2 Performance of CADS for various aperture sizes
CADS offers a better trade-off between depth and AIF quality compared to a naive dual-pixel sensor for various aperture sizes as shown in Fig. 1. To establish this, we evaluate the performance of CADS for various aperture sizes ranging from $f/4$ to $f/10$. We test this out on the same FlyingThings3D scenes. Apart from comparing CADS against naive dual-pixel sensing (Naive DP), we also compare it against coded and no-code variants of standard pixel sensor, referred to as Coded and Naive SP respectively. Table 2 shows depth estimation and AIF deblurring results on our simulated FlyingThings3D data, where the imaging system parameters are the same as mentioned before, with objects being placed in the [32 cm, 53 cm] range. With decreasing aperture size (increasing f-number), AIF prediction accuracy increases while depth prediction gets worse. Our proposed E2E CADS system consistently shows the best performance across all apertures and improves the defocus-disparity trade-off.
@@ 4.3 Comparison with existing dual-pixel works
We compare the performance of CADS with existing learning-based dual-pixel sensing works, namely DPDNet[2], DDDNet[21], Xin et al. [33], Punnapurath et al. [23] and Kim et al. [16]. Since existing works were developed for naive (no-code) dual-pixel sensors, we use the naive DP PSF blur to simulate captures for evaluation. For evaluating CADS, we render using the coded-aperture DP PSF. We use a validation subset of 2k images from the FlyingThings3D dataset to evaluate. The results are shown in Table 3. Most existing works estimate disparity which is related to defocus map by an unknown scale. For evaluation, we use the affine invariant version of MAE (AI(1)) for disparity estimation quality [9] and PSNR for AIF quality [^2] See Supplementary for metric definitions. Note that we use the normalized defocus map output by CADNet for this comparison instead of converting it to a depth map. Works that were designed for unidirectional disparity [21, 33] show poor results when tested on simulated scenes with bidirectional disparity between the DP images. Our proposed CADS outperforms existing methods on the simulated FlyingThings3D dataset. Moreover, CADNet trained on naive DP blurs, referred to as Naive DP in Table 3, also outperforms existing methods. We also perform a comparison on rendered images from NYUv2 dataset[28] and report the performance in the supplementary.
@@
Table 3: Comparison with existing DP methods. CADS offers the best AIF and disparity estimation quality on our simulated dataset. Red highlights best, orange highlights second best.